# Ocean Diviner: A Diffusion-Augmented Reinforcement Learning for AUV Robust Control in the Underwater Tasks

### v1

@@ IV Conclusion
This paper proposed a diffusion-augmented RL method for robust AUV control, overcoming key challenges in underwater trajectory planning and dynamic environment adaptation. The approach combines: (1) a diffusion-based trajectory generation framework with high-dimensional state encoding for improved long-horizon planning, (2) a hybrid learning architecture that merges diffusion-guided exploration with RL optimization for efficient and stable policy learning. Extensive simulations demonstrating superior robustness over conventional methods in complex marine environments.
Future work will focus on real-world validation to further verify the proposed method’s practicality and performance in the dynamic ocean environment.

### v2

@@ IV CONCLUSIONS
In this paper, we propose a diffusion-augmented RL framework for robust AUV control, addressing the challenges of nonlinear dynamics and uncertain underwater disturbances. The framework integrates: (1) a diffusion-based action generation mechanism with high-dimensional state encoding, enabling physically feasible actions and improved long-horizon planning, and (2) a hybrid learning architecture that combines diffusion-guided exploration with RL optimization, achieving sample-efficient and stable policy learning. Extensive simulations confirm superior robustness and flexibility compared to conventional control methods under dynamic marine conditions. Future work will focus on real-world experiments, further validating the adaptability of this framework for practical underwater operations.